The average length of time for students to register in the 1st semester at a certain school has been 50mins. A new registration procedure is being tested. If random samples of 20 students have an average of 45mins with a standard deviation of 9mins under the system, can you conclude that the new system is faster than the old? Use 0.01 level of significance.
1
Expert's answer
2020-05-25T16:29:47-0400

"a_0=50\\
n=20\\
\overline{x}=45\\
s=9\\
\alpha=0.01\\
H_0: \mu=a_0=50, H_1: \mu<a_0=50\\
\mu\text{ --- the average length of time for students}\\
\text{to register under the new system}.\\
\text{We assume that the length of time for students}\\
\text{to register has normal distribution}.\\
\text{We will use the following random variable as a criterion:}\\
T=\frac{(\overline{X}-a_0)\sqrt{n}}{s}\\
T\text{ has t-distribution with } k=n−1 \text{ degrees of freedom.}\\
\text{Observed value:}\\
t_{obs}=\frac{(45-50)\sqrt{20}}{9}\approx -2.48\\
\text{Critical value (one-sided):}\\
t_{cr}=t_{cr}(\alpha;k)=t_{cr}(0.01;19)\approx 2.54\\
(-\infty, -2.54)\text{ --- critical region}\\
t_{obs}\text{ does not fall into the critical region. So we accept } H_0.\\
\text{We cannot conclude that the new system is faster than the old}."